# PRISM [Uncaptioned image] of Opinions: A Persona‑Reasoned Multimodal Framework for User-centric Conversational Stance Detection

### v1

@@ 5.3 Main Results
In-Target Stance Detection. We report the experimental results on the U-MStance dataset under the in-target setup, as shown in Table 2, where the training and testing sets share identical targets. The results reveal several key observations. Notably, our proposed PRISM framework achieves the highest overall performance with an F1-avg of 68.49%, which substantially outperforms multimodal GPT4-1 with 66.24% and the text-only GPT-4 with 60.74%. This improvement highlights PRISM’s effectiveness and efficiency in capturing multimodal stance representations even with a relatively lightweight backbone. The results also confirm the importance of visual signals for stance understanding, as evidenced by consistent gains from multimodal models such as BERT-ViT compared to BERT, and GPT4-1 compared to GPT-4. However, social media image often convey subtle or context-dependent meanings that are challenging to interpret from surface-level visual features. By incorporating rationalized cross-modal grounding, PRISM effectively aligns textual and visual semantics, producing more accurate and context-aware stance predictions. Interestingly, the large models in our experiments generally underperform compared to smaller, specialized models. This may be because, despite their strong general language understanding, they are less effective at capturing fine-grained, conversation-specific, and multimodal cues that are critical for stance detection. In contrast, our PRISM leverages an LLM as its backbone, but through the integration of user persona and multimodality modeling, it effectively captures these nuanced cues, resulting in superior performance.
Cross-Target Stance Detection. We further evaluate generalization under the cross-target setting, where targets in the test set are unseen during training, as summarized in Table 3. Specifically, we conduct cross-target evaluations between related pairs such as Trump-Biden, Tesla-BMW, and Costco-Bitcoin, training on one and testing on the other. This scenario is more challenging, requiring the model to transfer stance understanding across targets with differing semantics and discourse style. LLMs are evaluated directly without additional fine-tuning, yielding performance identical to the in-target setup. We can observe that most smaller models suffer a notable degradation in performance, revealing their limited capacity to generalize beyond the specific targets encountered during training. Remarkably, our proposed PRISM framework maintains a relatively high and stable performance across all cross-target evaluations, demonstrating superior robustness and adaptability. This advantage stems from two components: the integration of user persona modeling, which allows the model to capture consistent user-level stance tendencies across topics, and the mutual task reinforcement mechanism, which jointly optimizes stance prediction and rationalized cross-modal grounding. Together, these components enable PRISM to establish a more target-invariant representation of stance, thereby enhancing its generalization capability across heterogeneous contexts.
@@ 5.4 Impact of Different Backbone
We further assess the adaptability of PRISM by integrating it with different MLLM backbones under both the in-target and cross-target settings, as illustrated in Table 2 and Table 3. Specifically, we replace the original Qwen-VL backbone with LLaVA and MiMo to examine how backbone choice affects performance. Across both evaluation setups, MiMo provides consistent performance gains across both evaluation setup. This finding demonstrates not only the effectiveness but also the generalizability of our framework, showing that PRISM can seamlessly adapt to stronger multimodal backbones while preserving its core advantages. The gains from MiMo likely stem from its enhanced multimodal reasoning capabilities and more robust cross-domain representation learning. These properties align well with PRISM’s design, enabling more precise modeling of nuanced stance cues and improving the transferability of stance understanding across heterogeneous targets.
Table 4: Experimental results (%) of the ablation study on in-target and cross-target multimodal conversational stance detection. w/o means without.
@@ 5.5 Ablation Study
To evaluate the contribution of each component in our PRISM framework, we conduct a comprehensive ablation study with the results presented in Table 4. Removing the persona representation $p_{u_{N}}$ (w/o Persona) yields a consistent performance drop of approximately 2%-3%, underscoring the significance of modeling user-specific characteristics. This component enables the framework to capture individualized traits, thereby enhancing its ability to accurately interpret user-centric stance expressions across conversational contexts. Eliminating the intent-aware caption $\hat{x}$ (w/o Intent) results in a more substantial degradation, highlighting the crucial role of the rationalized cross-modal grounding module. By explicitly aligning the pragmatic intent of images with the accompanying textual context, this component enables the model to reason about subtle or implicit visual cues, which are often critical for correctly inferring stance in multimodal conversations. Lastly, discarding the mutual task reinforcement mechanism and reverting to a single-task stance detection objective (w/o Mutual) also leads to a noticeable decline in performance. This confirms that jointly optimizing stance detection with auxiliary generation task encourages beneficial interactions between multimodal understanding and stance reasoning, ultimately producing more robust and transferable representations.
@@ 5.6 Case Study
To visually demonstrate the superiority of our method and the inherent challenges of the task, we conduct a case study, as illustrated in the figure 5. As shown in Case (a), both GPT4-1 and LLaVA fail to detect the sarcastic implication conveyed jointly by the text and image, incorrectly assigning a neutral stance. These models tend to rely on surface-level cues and struggle with recognizing pragmatic intent. In contrast, our PRISM accurately predicts an Against stance. This improvement stems from its rationalized cross-modal grounding, which captures the user’s implicit criticism expressed through the comparison embedded in the image, as well as its persona-aware interpretation that contextualizes the user’s communicative tendencies.
Meanwhile, Case (b) presents an even more subtle scenario, where sarcasm is intertwined with the user’s disposition. Again, GPT4-1 and LLaVA misclassify the stance as neutral, likely due to their lack of user-level modeling and reliance solely on the current conversational snippet. PRISM, however, leverages its persona module to interpret the user’s longitudinal behavioral patterns. As shown in the PRISM persona visualization, the user exhibits a prominently high Neuroticism score, reflecting a more emotionally expressive and critical communication style. Incorporating such personality cues enables PRISM to infer the user’s consistent critical orientation, allowing it to correctly identify the Against stance even in contexts where the sarcastic intent is subtle and easily overlooked.
@@ 5.7 Impact of Conversation Depth
We further investigate how conversational complexity influences stance detection by evaluating PRISM across varying conversation depths, as shown in Figure 6. Overall, performance shows a gradual decline as depth increases, which is expected since deeper conversations involve more intricate discourse structures, longer-range dependencies, and greater contextual ambiguity, making the final stance harder to infer. Notably, for the targets Costco and Bitcoin, PRISM maintains relatively stable performance across depths. This robustness is largely attributed to the rationalized cross-modal grounding module, whose intent-aware image captions help align visual and textual cues and preserve salient multimodal information, thereby mitigating the challenges posed by longer conversational contexts.
@@ 6 Conclusion
This paper presents two critical limitations in MCSD, including pseudo-multimodality, where visual cues are limited to source posts, and user homogeneity, which neglects individual differences. To address these, we propose U-MStance, the first user-centric MCSD dataset with multimodal and user information. Furthermore, we design PRISM, a framework that distills stable user personas, grounds image intent, and reinforces stance understanding through multitask learning. Entensive Experiments on U-MStance demonstrate that PRISM consistently outperforms strong baselines and exhibits robust generalization.

### v2

@@ 5.3 Main Results
In-Target Stance Detection. We report the experimental results on the U-MStance dataset under the in-target setup, where the training and testing sets share identical targets. As shown in Table 2, our proposed PRISM framework achieves the highest overall performance with an F1-avg of 68.49%, which substantially outperforms multimodal GPT4-1 with 66.24% and the text-only GPT-4 with 60.74%. This improvement highlights PRISM’s effectiveness and efficiency in capturing multimodal stance representations even with a relatively lightweight backbone. The results also confirm the importance of visual signals for stance understanding, as evidenced by consistent gains from multimodal models such as BERT-ViT compared to BERT, and GPT4-1 compared to GPT-4. By incorporating rationalized cross-modal grounding, PRISM effectively aligns textual and visual semantics, producing more accurate and context-aware stance predictions. Interestingly, the large models, such as LLAVA and MiMo, generally underperform compared to smaller, specialized models like TMPT. Despite their strong general language understanding, these larger models may be less effective at capturing fine-grained, conversation-specific, and multimodal cues that are critical for stance detection. In contrast, our PRISM leverages user persona and multimodal modeling to effectively capture nuanced cues, enabling its LLM backbone to achieve superior performance.
Cross-Target Stance Detection. We further evaluate generalization under the cross-target setting, where targets in the test set are unseen during training, as summarized in Table 3. Specifically, we conduct cross-target evaluations between pairs such as Trump-Biden, Tesla-BMW, and Costco-Bitcoin, training on one and testing on the other. This scenario is more challenging, requiring the model to transfer stance understanding across targets with differing semantics and discourse style. We can observe that most smaller models suffer a notable degradation in performance, revealing their limited capacity to generalize beyond the specific targets encountered during training. Remarkably, our proposed PRISM framework maintains a relatively high and stable performance across all cross-target evaluations, demonstrating superior robustness and adaptability. This advantage stems from two components: the integration of user persona modeling, which allows the model to capture consistent user-level stance tendencies across topics, and the mutual task reinforcement mechanism, which jointly optimizes stance prediction and rationalized cross-modal grounding. Together, these components enable PRISM to establish a more target-invariant representation of stance, thereby enhancing its generalization capability across heterogeneous contexts.
@@ 5.4 Impact of Different Backbone
We further assess the adaptability of PRISM by integrating it with different MLLM backbones under both the in-target and cross-target settings, as illustrated in Table 2 and Table 3. Specifically, we replace the original Qwen-VL backbone with LLaVA and MiMo to examine how backbone choice affects performance. Across both evaluation setups, MiMo provides consistent performance gains across both evaluation setups. This finding demonstrates not only the effectiveness but also the generalizability of our framework, showing that PRISM can seamlessly adapt to stronger multimodal backbones while preserving its core advantages. The gains from MiMo likely stem from its enhanced multimodal reasoning capabilities and more robust cross-domain representation learning. These properties align well with PRISM’s design, enabling more precise modeling of nuanced stance cues and improving the transferability of stance understanding across heterogeneous targets.
Table 4: Experimental results (%) of the ablation study. w/o means without.
@@ 5.5 Ablation Study
To evaluate the contribution of each component in our PRISM framework, we conduct a comprehensive ablation study with the results presented in Table 4. Replacing the persona representation $p_{u_{N}}$ with the user’s original raw information (w/o Persona) consistently drops performance across all targets. This underscores the critical role of persona distillation in filtering inherent noise and extracting task-relevant characteristics, enabling PRISM to more precisely interpret individualized, user-centric stance expressions.
Substituting intent-aware captions $\hat{x}$ with raw visual features (w/o Intent) results in a more substantial degradation, highlighting the crucial role of rationalized cross-modal grounding. By explicitly aligning the image pragmatic intent with textual context, this component enables the PRISM to correctly infer subtle or implicit visual cues essential for multimodal stance detection. Lastly, discarding the mutual task reinforcement mechanism and reverting to a single-task stance detection objective (w/o Mutual) also leads to a noticeable decline in performance. This confirms that jointly optimizing stance detection with auxiliary generation task encourages beneficial interactions between multimodal understanding and stance reasoning, ultimately producing more robust and transferable representations.
@@ 5.6 Case Study
To visually demonstrate the superiority of our method and the inherent challenges of the task, Figure 4 provides a case study. In Case (a), both GPT4-1 and LLaVA fail to detect the sarcastic implication conveyed by joint multimodal cues, incorrectly assigning a neutral stance due to their reliance on surface-level information. In contrast, our PRISM accurately predicts an Against stance. This improvement stems from its rationalized cross-modal grounding, which captures the user’s implicit criticism expressed through the comparison embedded in the image, as well as its persona-aware interpretation of the user’s communicative tendencies.
Meanwhile, Case (b) presents a more subtle scenario, where sarcasm is intertwined with the user’s disposition. Again, GPT4-1 and LLaVA misclassify the stance, likely due to their lack of user-level modeling and reliance solely on the current conversational snippet. PRISM, however, correctly identifies the Against stance by leveraging its persona module to interpret longitudinal behavioral patterns. Specifically, the persona visualization highlights a high Neuroticism score, reflecting an emotionally expressive and critical style. Incorporating such personality cues enables PRISM to infer the user’s consistent critical orientation even when sarcasm is subtle.
@@ 5.7 Impact of Conversation Depth
We further investigate evaluate how PRISM’s stance detection performance is influenced by varying conversational depths. As shown in Figure 5, performance shows a gradual decline as depth increases, which is expected since deeper conversations involve more intricate discourse structures, longer-range dependencies, and greater contextual ambiguity, making the final stance harder to infer. Notably, for the targets Costco and Bitcoin, PRISM maintains relatively stable performance across depths. This robustness is largely attributed to the rationalized cross-modal grounding module, whose intent-aware image captions help align visual and textual cues and preserve salient multimodal information, thereby mitigating the challenges posed by longer conversational contexts.
@@ 6 Conclusion
This paper presents two critical limitations in MCSD, including pseudo-multimodality, where visual cues are limited to source posts, and user homogeneity, which neglects individual differences. To address these, we propose U-MStance, the first user-centric MCSD dataset with multimodal and user information. Furthermore, we design PRISM, a framework that distills stable user personas, grounds image intent, and reinforces stance understanding through multitask learning. Extensive Experiments on U-MStance demonstrate that PRISM consistently outperforms strong baselines and exhibits robust generalization.
Limitations
Despite the significant improvements achieved by our model, certain limitations remain. First, while U-MStance spans diverse domains including politics, automotive, consumer goods, and cryptocurrency, it lacks coverage of highly specialized or professional sectors, such as specific legal statutes or cutting-edge scientific controversies. Second, as illustrated in Figure 5, model performance exhibits a declining trend as conversational depth increases. In extremely deep or multi-branch discussion threads, the model still encounters challenges in perfectly capturing subtle logical transitions and long-range contextual drifts.
Appendix A Detailed Dataset Statistics
Figure 6 illustrates the distribution of instances across different labels. Table 5 illustrates the distribution of instances across different conversational depths, including the number of users and images for each target. Within our collected dataset, 36% of the posts contain a single image, while 63% include two or more images, which means visual information is also present within the comments. In general, our dataset features more extensive conversational depth compared to existing conversational stance detection datasets, and achieves 100% vision-related coverage.
Figure 6: The distribution of stance categories in our U-MStance.
Table 5: Comparison of Depth and Interaction Data Across Different Entities
Depth Avg. WC Trump Biden Tesla BMW Costco Bitcoin Total
User 5087 3371 3896 3956 4519 4123 24952
Image 47 40 55 102 20 21 285
Post 1 17.96 16 27 14 19 8 13 97 (0.24%)
Comment 2 27.38 927 1011 723 741 768 1186 5356 (13.39%)
3 23.71 2306 2096 1982 1816 2288 2647 13135 (32.84%)
4 27.65 1645 1189 1402 1176 1470 1292 8174 (20.43%)
5 29.99 1076 694 928 821 1001 846 5366 (13.41%)
6 31.52 795 427 621 637 541 517 3538 (8.84%)
7 35.98 529 231 430 385 316 304 2195 (5.49%)
8 39.59 333 106 264 290 168 173 1334 (3.33%)
9 45.02 231 65 150 148 108 106 808 (2.02%)
Appendix B Model Details
Table 6: Model details including parameters, access type, and their respective sources.
Model Parameters Access Source
BERT ¡1B Local Google
RoBERTa ¡1B Local Meta
PoliBERTweet ¡1B Local Kawintiranon et al.
LLaMA-2-7B 7B Local Meta
GPT-3.5 Unknown API OpenAI
GPT-4 Unknown API OpenAI
BERT-ViT ¡1B Local Liang et al.
TMPT ¡1B Local Liang et al.
LLaVA-1.5-7B 7B Local Microsoft
Qwen2.5-VL-7B 7B Local Alibaba
MiMo-VL-7B 7B Local Xiaomi
GPT4-1 Unknown API OpenAI
Appendix C Prompt Templates
To support the reproducibility and interpretability of our methodology and experimental settings, we provide the comprehensive set of prompt templates used in this work. Specifically, Table 7 details the prompts for User Persona Distillation; Table 8 presents the two-stage prompt protocol for Image Intent Extraction; and Table 9 outlines the prompt employed for the baseline models.
B.1 User Persona Distillation System Prompt: You are an expert in computational psychology and linguistic analysis. Your task is to analyze a user’s Big Five personality traits (Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism) based on their historical social media activity. User Prompt: ### Task Description: Analyze the provided user profiles, including their posts and comments. For each of the five traits, provide a qualitative description of how the user exhibits this trait and a quantitative score on a scale of 1-5 (1: Very Low, 5: Very High). ### Big Five Traits Definition: 1. Openness to Experience: Curiosity, originality, and openness to new ideas. 2. Conscientiousness: Organization, responsibility, and goal-oriented behavior. 3. Extraversion: Sociability, assertiveness, and emotional expressiveness. 4. Agreeableness: Trust, altruism, kindness, and affection. 5. Neuroticism: Emotional instability, anxiety, and irritability. ### Input Data: User ID: {{user id}} Recent Posts: {{post list}} Recent Comments: {{comment list}} Output:
Table 7: Prompt template for User Persona Distillation. The model outputs a summary of Big Five personality traits for users.
B.2 Image Intent Extraction Step 1: System Prompt: You are an expert image analyzer. User Prompt: ### Task Description: You will be given an image. Your task is to produce a detailed and structured description of the image. Please follow the required steps: 1. **Overall Scene Description** Describe the setting, environment, and general context of the image. 2. **Key Objects and Entities** Identify and describe the notable objects, people, or elements in the image, including their appearance, attributes, and spatial relationships. 3. **Fine-Grained Visual Details** Highlight colors, expressions, textures, gestures, body language, or any small details that may convey meaning. 4. **Style and Emotional Tone (if applicable)** Describe the mood, atmosphere, artistic style, or emotional impression conveyed by the image. ### Input: ¡image¿ ### Output Requirment: Make sure your description is precise, neutral, and avoids assumptions unless clearly supported by visual evidence. Directly output the description text**, starting with **‘Image Description:‘** followed by your concise answer. Output: Step 2: System Prompt: You are an expert in multimodal conversation analysis. User Prompt: ### Task Description: Determine the **intention or communicative purpose** of the user who posted the image in the conversation. Consider the social and semantic context of the conversation, as well as the content of the image. You are given: 1. A **description of the image**, which has been generated by a visual-language model. 2. A **conversation history** between multiple participants. - In this history, the placeholder ”¡image¿” marks where the image was posted by the user. If the conversation contains multiple images, carefully identify the **index or order** of the images (e.g., ¡image1¿, ¡image2¿, ¡image3¿…) and determine where the **current target image** appears in the conversation before analyzing it. ### Requirment: Follow these steps carefully: 1. Review the conversation history and understand the flow of dialogue. 2. Interpret the meaning of the image based on the provided description. 3. Analyze how the image contributes to the conversation — for example, whether it expresses emotion, humor, evidence, sarcasm, agreement, mockery, or provides visual information. 4. Infer the **most likely intention** of the user when posting the image. 5. Directly output the intent text**, starting with **‘Intent:‘** followed by your concise answer. ### Image Description: The image captures a bustling urban scene at night, likely Times Square in New York City, characterized by its iconic billboards and vibrant advertisements. The foreground features … ### Conversation History: User youyouxue: We just drove from San Francisco to New York City in a Tesla Model X without using a single drop of gas! ¡image1¿ User salec1:I did a shorter version of the same thing (Miami to NYC) last month. It’s all about … User KickAClay:I always wondered if any Tesla, after inputting your route … User timdorr:Yes, it does exactly this. Set your destination and it will find Superchargers along your route, tell you how long to … User chupacabrando:How much longer does it take … Output:
Table 8: Prompt template for Image Intent Extraction. The model outputs a detailed and structured description of the image.
B.3 Zero-shot Stance Detection System Prompt: You are an expert linguist specializing in conversational analysis and social media stance detection. User Prompt: ### Task Description: Analyze the provided conversation flow and determine the stance of the current speaker toward the specific Target listed below. The attached image(s) correspond to the sequential labels (e.g., ¡image1¿, ¡image2¿) mentioned in the text. ### Target: Tesla ### Stance Classes: [’favor’,’against’,’none’] ### Conversation History: User youyouxue: We just drove from San Francisco to New York City in a Tesla Model X without using a single drop of gas! ¡image1¿ User salec1:I did a shorter version of the same thing (Miami to NYC) last month. It’s all about … User KickAClay:I always wondered if any Tesla, after inputting your route … User timdorr:Yes, it does exactly this. Set your destination and it will find Superchargers along your route, tell you how long to … User chupacabrando:How much longer does it take … ### Current Speaker’s Reddit Profile: Active Communities: — Historical Posts — — Historical Comments — ### Output Format (JSON): s̈tance:̈ ”favor” — ”against” — ”none” Output:
Table 9: Prompt template for Multimodal Conversational Stance Detection. The model outputs a stance class for the current instance.